Calculate the MINIMUM value of the suitable range of fluid required for a 37 year old male patient weighing 240 pounds and 6 feet tall. This patient has suffered from severe trauma.
Answer #1

The Holliday - Segar 4-2-1 Rule is used to estimate the maintainance fluid requirement .

Weight of the patient is 240 pound = 0.45×240 = 108 kg

Height of the patient is 6feet

Applying the formula :-

# For more than 20kg weight, daily fluid requirement is

= 1500 ml + 20ml / kg/ day for every kg > 20

= 1500 ml+ 20 ml× 108kg

= 1500 ml + 2160 ml

= 3660ml of fluid is required in 24hours

# For more than 20kg weight, hourly fluid requirement is

Weight the patient = 108kg

= First 10 kg is 4ml/kg = 40ml

= Second 10kg is 2ml/kg = 20ml

Then 1 ml/kg for rest of the weight = 88ml

Total = 40 +20+ 88 = 148 ml /hr .